if he has 3 times as many as her, they have a total of 36 together. how many does she have?
Answer by lmeeks54(111)

Let b = # the boy has
Let g = # the girl has
...
we have enough information to generate two equations for the two unknowns:
...
He has 3 times as many as her:
b = 3 * g ... or b = 3g
...
Together, they have 36:
b + g = 36
...
With two equations and two unknowns, if we have one of the equations cast in terms of one of the unknowns, we can substitute that equality into the other equation so we have one equation with one unknown (this is a very useful general method for solving lots of problems like this):
...
take, b = 3g and insert it into b + g = 36:
...
3g + g = 36
4g = 36
g = 36/4
g = 9
...
Take g = 9 and insert it back into either of the original equations to solve for b:
...
b = 3g
b = 3(9)
b = 27
...
Always check your work... We do that by going back to the original equations with the computed values to see if both equations work:
...
b + g = 36
27 + 9 = 36
36 = 36 ... checks
...
cheers,
Lee